Bond dissociation energy of AB molecules is 300kj/mol.the number of moles of photons other wavelength 6625 Angstrom. Requires to dissociate 3molesof AB Molecules is

Grade:12th pass

1 Answers

Arun
5 years ago
First convert 300kj/mol to joules ie
(300*10^3)/(6.022*10^23)= 0.498*10^-18------[1]

Now we need to find the no of moles photons for 3 moles of AB molecule, so
3*E=nhc/λ {where n=no of moles of photons; E = [1]= 0.498*10^-18 and λ = 6625*10^-10 m}
=> 3*0.498*10^-18=(n*6.625*10^-34*3*10^8)/6625*10^-10
Solving the above equation we obtain
n = 4.98 ≈5 mol